OpenStudy (anonymous):

If ABC DEF and the scale factor from abc to def is 3/4 what is the length of df? (ba=12)(ac=15)

OpenStudy (mandre):

Calculate BC using pythagoras then you multiply the answer by 3/4 or 0.75

OpenStudy (anonymous):

lol what is a pythagora

OpenStudy (anonymous):

i didnt learn anything about that.

OpenStudy (mandre):

Nope DF = AC * (3/4)

OpenStudy (mandre):

Since they are the equivalent sides. No pythagoras needed.

OpenStudy (mandre):

Draw the triangles next to each other then you will see.

OpenStudy (anonymous):

okay, so what is df?

OpenStudy (mandre):

Just make sure the equivalent sides are in the same place. 

OpenStudy (mandre):

What is ac?

OpenStudy (anonymous):

so just make df 15

OpenStudy (anonymous):

thanks :)!

OpenStudy (mandre):

you forgot the scale of 3/4

OpenStudy (anonymous):

howd you draw that.

OpenStudy (anonymous):

oh idk what a scale is or how to use that.

OpenStudy (mandre):

DF = AC * (3/4) Take AC's length then multiply by (3/4)

OpenStudy (anonymous):

how would i multiply that?

OpenStudy (mandre):

Take 15 x 3 Take that answer then divide it by 4.

OpenStudy (anonymous):

i got 11.25

OpenStudy (mandre):

Yup.

OpenStudy (anonymous):

thanks so much, can you help me with anothe problem

OpenStudy (mandre):

Lol. ok

OpenStudy (anonymous):

if abc~def what is the length of ef ( im about to try to draw it out)

OpenStudy (anonymous):



OpenStudy (mandre):

You get the scale by dividing 2 equivalent sides.

OpenStudy (mandre):

Scale = DF / AC = 10 / 15

OpenStudy (mandre):

Now EF = BC * Scale

OpenStudy (mandre):

You can also use DE / AB = 4 / 6 It will be the same 4 / 6 = 10 / 15

OpenStudy (anonymous):

lol im so lost now

OpenStudy (mandre):

What is 10 / 15?

OpenStudy (mandre):

You can even use 10 / 15 as your scale

OpenStudy (mandre):

EF = BC * Scale So take BC then multiply it by 10. Then you divide that answer by 15.

OpenStudy (mandre):

BC = 12 What is 12 x 10?

OpenStudy (anonymous):

ef=8

OpenStudy (mandre):

There you go

OpenStudy (anonymous):

thanks, now another problem lol?

OpenStudy (mandre):

Noooo. Post it as a new question. I've gotta get back to work. :)